# Nuclear Toeplitz operators between Fock spaces

### v1

@@
Abstract.
We study Toeplitz operators with measure-valued symbols acting between Fock spaces. Given $1\leq p,q\leq\infty$ and a Borel measure $\mu$ on $\mathbb{C}$, we investigate when the associated Toeplitz operator
$T_{\mu}:F^{p}_{\alpha}\to F^{q}_{\alpha}$
belongs to the nuclear class. For positive measures $\mu$ and in the range $1\leq q\leq p\leq\infty$, we obtain necessary and sufficient conditions for the nuclearity of $T_{\mu}$ in terms of the Berezin transform of $\mu$. As a consequence, nuclearity in this setting exhibits a rigidity property: if $T_{\mu}$ is nuclear from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $q\leq p$, then it is nuclear for all such $q$. In the case $p<q$, we show that the situation is more delicate. We provide separate necessary and sufficient conditions for nuclearity, indicating that the Berezin transform alone does not yield a complete characterization. The proofs rely on tools from Banach space operator theory combined with kernel estimates on Fock spaces. Our results extend naturally to Fock spaces on $\mathbb{C}^{n}$.
Key words and phrases:
@@ Abstract.
1991 Mathematics Subject Classification:
Primary 47B10 Secondary 47B35
*Corresponding author.
1. Introduction
@@
Let $\alpha>0$ and $p>0$, the notation $L_{\alpha}^{p}$ refers to the space of all Lebesgue measurable functions $f$ on $\mathbb{C}$ for which $f(z)\mathrm{e}^{-\alpha|z|^{2}/2}\in L^{p}(\mathbb{C},\mathrm{d}A)$, where $\mathrm{d}A$ is the Lebesgue area measure on $\mathbb{C}$. For $f\in L_{\alpha}^{p}$, its norm is given by
$\|f\|_{\alpha,p}^{p}=\frac{p\alpha}{2\pi}\int_{\mathbb{C}}\left|f(z)\mathrm{e}^{-\frac{\alpha}{2}|z|^{2}}\right|^{p}\,dA(z).$
@@
Correspondingly, for $\alpha>0$ and $p=\infty$, we denote by $L_{\alpha}^{\infty}$ the space of measurable functions $f$ such that
$\|f\|_{\infty,\alpha}=\operatorname{ess\,sup}\left\{|f(z)|\mathrm{e}^{-\alpha|z|^{2}/2}:z\in\mathbb{C}\right\}<\infty.$
@@
The Fock space $F_{\alpha}^{p}$ is defined as the subspace of $L_{\alpha}^{p}$ consisting of entire functions. It is straightforward to verify that $F_{\alpha}^{p}$ becomes a Banach space when $1\leq p\leq\infty$. Let $f_{\alpha}^{\infty}$ be the closed subspace of $F_{\alpha}^{\infty}$ consisting of $f$ satisfying
$\lim_{|z|\to\infty}f(z)e^{-\frac{\alpha}{2}|z|^{2}}=0.$
@@
When $p=2$, the space $F_{\alpha}^{2}$ is a Hilbert space equipped with the inner product
$\langle f,g\rangle_{F_{\alpha}^{2}}=\frac{\alpha}{\pi}\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\,dA(z).$
@@
As a reproducing kernel Hilbert space, its reproducing kernel is $K_{\alpha}(z,w)=e^{\alpha z\overline{w}}$. The normalized reproducing kernel at $z$ is denoted by $k_{z}(w)=e^{\alpha\overline{z}w-\frac{\alpha}{2}|z|^{2}}$, which is a unit vector in $F_{\alpha}^{p}$ for every $0<p\leq\infty$.
For $1\leq p<\infty$, let $p^{\prime}$ be the conjugate exponent satisfying $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. We can identify the dual of $F_{\alpha}^{p}$ with $F_{\alpha}^{p^{\prime}}$, and the dual of $f_{\alpha}^{\infty}$ with $F_{\alpha}^{1}$, via the integral pairing
$\langle f,g\rangle_{F_{\alpha}^{2}}=\frac{\alpha}{\pi}\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\,dA(z),$
and under this identification, the norms on the dual spaces and the Fock spaces are equivalent (see [19, Corollary 2.25, Theorem 2.26]).
@@
$d\lambda_{\alpha}(z)=\frac{\alpha}{\pi}e^{-\alpha|z|^{2}}\,dA(z).$
Define the integral operator $P_{\alpha}$ on $L_{\alpha}^{p}$ by
$P_{\alpha}f(z)=\int_{\mathbb{C}}K_{\alpha}(z,w)f(w)\,d\lambda_{\alpha}(w),\quad f\in L_{\alpha}^{p}.$
It is known that $P_{\alpha}:L_{\alpha}^{p}\to F_{\alpha}^{p}$ is bounded for all $p$, and it acts as the identity on $F_{\alpha}^{p}$, that is, $P_{\alpha}f=f$ for $f\in F_{\alpha}^{p}$ (see [19, Corollary 2.22]).
@@
$T_{\varphi}(f)=P_{\alpha}(\varphi f),\quad f\in F_{\alpha}^{2}.$
More generally, if $\mu$ is a complex Borel measure on $\mathbb{C}$, the Toeplitz operator $T_{\mu}$ is given by
$T_{\mu}(f)(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}K_{\alpha}(z,w)f(w)\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w),\quad z\in\mathbb{C},$
provided that the integrability condition
$\int_{\mathbb{C}}|K_{\alpha}(z,w)|\mathrm{e}^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty$ (1.1)
holds for every $z\in\mathbb{C}$. Due to the exponential nature of the kernel, condition (1.1) is equivalent to
$\int_{\mathbb{C}}|K_{\alpha}(z,w)|^{2}\mathrm{e}^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty,\quad z\in\mathbb{C}.$ (1.2)
Let
@@
The set $K$ is dense in $F_{\alpha}^{p}$ for every $1\leq p<\infty$ and also in $f_{\alpha}^{\infty}$. Under assumption (1.2), the operator $T_{\mu}$ is well-defined on $K$ by the Cauchy-Schwarz inequality, and hence on a dense subset of each $F_{\alpha}^{p}$ ($1\leq p<\infty$) and $f_{\alpha}^{\infty}$. For a measure $\mu$ satisfying (1.1), its Berezin transform is defined as
@@
Toeplitz operators on Fock spaces form an important class of operators connecting complex analysis, operator theory, and the theory of Banach function spaces. Owing to the reproducing kernel structure of Fock spaces, Toeplitz operators provide a natural framework for studying how analytic and measure-theoretic properties of symbols are reflected in operator-theoretic behavior. The theory of Toeplitz operators on Fock spaces has been extensively developed in the Hilbert space setting. For the classical Fock space $F^{2}_{\alpha}$, Toeplitz operators with bounded or measure-valued symbols have been studied from various perspectives, including boundedness, compactness, and Schatten class membership, see references [1, 2, 3, 9]. In particular, Isralowitz and Zhu [13] obtained complete characterizations of boundedness, compactness, and Schatten class properties of Toeplitz operators with positive Borel measures as symbols.
Beyond the Hilbertian framework, Hu and Lv [11, 12] systematically investigated Toeplitz operators acting between different Fock spaces $F^{p}_{\alpha}$ and $F^{q}_{\alpha}$ for $1<p,q<\infty$. They characterized boundedness and compactness of Toeplitz operators with positive Borel measure symbols in terms of averaging functions and Berezin-type transforms. A notable feature of their results is a rigidity phenomenon specific to Fock spaces: boundedness (or compactness) from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $p\leq q$ automatically implies boundedness (or compactness) for all such $q$. This behavior differs substantially from what is known in the Bergman space setting. The groundbreaking work of Hu and Lv has stimulated extensive scholarly investigations into Toeplitz and Hankel operators across diverse weighted Fock spaces [16, 19].
In contrast to the well-developed theory of bounded and compact Toeplitz operators, considerably less is known about Toeplitz operators belonging to finer operator ideals on Fock spaces, especially in the non-Hilbertian case. Among these ideals, the class of nuclear operators plays a distinguished role in Banach space operator theory. Nuclearity represents a strong form of compactness and depends sensitively on the geometry of the underlying Banach spaces. As a result, techniques based on Hilbert space methods or Schatten class theory are generally insufficient to address nuclearity questions when $p\neq 2$.
The aim of this paper is to study nuclear Toeplitz operators acting between Fock spaces. Given $1\leq p,q\leq\infty$ and a Borel measure $\mu$ on $\mathbb{C}$, we investigate conditions under which the Toeplitz operator
$T_{\mu}:F^{p}_{\alpha}\longrightarrow F^{q}_{\alpha}$
belongs to the nuclear class $N(F^{p}_{\alpha},F^{q}_{\alpha})$. Our focus is primarily on Toeplitz operators with positive symbols, although several results extend to complex measures under suitable assumptions.
Our main results provide a detailed description of nuclear Toeplitz operators in different parameter regimes. When $1\leq q\leq p\leq\infty$ and $\mu\geq 0$, we obtain necessary and sufficient conditions for $T_{\mu}$ to be nuclear in terms of the integrability of the Berezin transform of $\mu$ (Theorem 3.7). As a consequence, we show that nuclearity in this range exhibits a rigidity property analogous to that observed for boundedness: if $T_{\mu}$ is nuclear from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $q\leq p$, then it is nuclear for all such $q$.
When $p<q$, the situation becomes more subtle. In this case, we establish separate necessary and sufficient conditions for nuclearity, showing that the Berezin transform alone does not yield a complete characterization. This reflects an inherent asymmetry between the domain and target spaces in the nuclear setting and distinguishes the nuclear theory from previously studied mapping properties of Toeplitz operators on Fock spaces.
The proofs rely on tools from Banach space operator theory, including duality properties of nuclear operators, approximation and Radon–Nikodým properties of Fock spaces, and precise kernel estimates associated with the Berezin transform. The arguments presented here extend naturally to Fock spaces on $\mathbb{C}^{n}$, and the main results remain valid in higher dimensions without essential changes.
The paper is organized as follows. In Section 2, we recall basic definitions and preliminary results concerning nuclear operators and Fock spaces. Section 3 is devoted to the characterization of nuclear Toeplitz operators between Fock spaces and contains our main results. In Section 4, we prove density results showing that Toeplitz operators with compactly supported continuous symbols are dense in the space of nuclear operators between Fock spaces with respect to the nuclear norm. As a consequence, we recover and extend known results on trace ideals in the Hilbert space case $p=2$.
2. Preliminaries
This section collects known results and establishes basic facts concerning nuclear operators, which will be used in the subsequent sections.
@@ 2. Preliminaries
for sequences $(x_{j}^{\prime})\subset X^{\prime}$ and $(y_{j})\subset Y$ satisfying $\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y}<\infty$. Let $x^{\prime}\in X^{\prime},y\in Y$ , we use the notation $x^{\prime}\otimes y$ for the rank-one operator from $X$ to $Y$ defined by
$x^{\prime}\otimes y(x)=\langle x,x^{\prime}\rangle y,\quad x\in X.$
The nuclear norm of $T$ is given by
$\|T\|_{\mathcal{N}}:=\inf\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y},$
The infimum in this definition ranges over all representations of the type (2.1). we write $\mathcal{N}(X,Y)$ for the class of all nuclear operators mapping $X$ into $Y$. Endowed with the norm $\|\cdot\|_{\mathcal{N}}$, it becomes a Banach space, and the following estimate holds for any $A\in\mathcal{L}(X_{0},X)$, $T\in\mathcal{N}(X,Y)$, and $B\in\mathcal{L}(Y,Y_{0})$:
$\|BTA\|_{\mathcal{N}}\leq\|A\|\,\|T\|_{\mathcal{N}}\,\|B\|.$
In the case where $E$ and $F$ coincide, we write $\mathcal{N}(E)$ for $\mathcal{N}(E,E)$. We say that a Banach space $E$ has the approximation property provided that for any compact set $K\subset E$ and any $\epsilon>0$, one can find a finite-rank operator $L:E\to E$ satisfying
$\|x-Lx\|<\epsilon\quad\text{for all }x\in K.$
For a Banach space $E$ with the approximation property, the nuclear trace of an operator $T\in\mathcal{N}(E)$ can be defined by
$\operatorname{Tr}_{E}(T)=\sum_{j=1}^{\infty}\langle x_{j},x_{j}^{\prime}\rangle,$
where $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes x_{j}$ is any representation as in (2.1). It follows from [10] Theorem V.1.2 that the value of the trace does not depend on the particular representation chosen. Standard references for the general theory of operator ideals include [5, 6, 10, 17].
@@
$G(E)=\int_{E}g\,d\mu\qquad\text{holds for all }E\in\Sigma,$
we say $X$ has the Radon-Nikodým property.
For each nonnegative integer $n$, define
$e_{n}(z)=\sqrt{\frac{\alpha^{n}}{n!}}z^{n}.$
It is well known that $\{e_{n}\}_{n=0}^{\infty}$ forms an orthonormal basis for $F_{\alpha}^{2}$. For a general exponent $1<p<\infty$, the same system $\{e_{n}\}_{n=0}^{\infty}$ is also a Schauder basis for $F_{\alpha}^{p}$, as can be seen by expanding each $f\in F_{\alpha}^{p}$ in a Taylor series about the origin.
@@
For $1<p<\infty$, the Fock space $F_{\alpha}^{p}$ is a Banach space that enjoys both the approximation property and the Radon–Nikodým property.
The approximation property follows from the fact that $F_{\alpha}^{p}$ has a Schauder basis (see [9, Proposition 2.1]). The Radon–Nikodým property holds because $F_{\alpha}^{p}$ is reflexive (see [7, Corollary 3.13]). ∎
@@
$\displaystyle T^{*}(y^{\prime})(x)$ $\displaystyle=y^{\prime}(Tx)$
$\displaystyle=y^{\prime}(\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y_{j})$
$\displaystyle=\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y^{\prime}(y_{j})$
$\displaystyle=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}(y^{\prime})x_{j}^{\prime}(x).$
Then $T^{*}=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}\otimes x_{j}^{\prime}$, and $\|T^{*}\|_{\mathcal{N}}\leq\sum_{j=1}^{\infty}\|y_{j}^{\prime\prime}\|_{Y^{\prime\prime}}\|x_{j}^{\prime}\|_{X^{\prime}}=\sum_{j=1}^{\infty}\|{y_{j}\|_{Y}\|x_{j}^{\prime}\|_{X^{\prime}}}$, the inequality $\|T^{*}\|_{\mathcal{N}}\leq\|T\|_{\mathcal{N}}$ follows by taking the infimum over all representations of the form (2.1). On the other hand, let $K\in\mathcal{N}(Y^{\prime},X^{\prime})$, since $Y$ is reflexive, there are sequences $(a_{j}^{\prime})\in X^{\prime}$, $(b_{j})\in Y$ such that $K=\sum_{j=1}^{\infty}b_{j}^{\prime\prime}\otimes a_{j}^{\prime}$, let $L=\sum_{j=1}^{\infty}a_{j}^{\prime}\otimes b_{j}\in\mathcal{N}(X,Y)$, It is easy to check that $L^{*}=K$ and $\|L\|_{\mathcal{N}}\leq\|K\|_{\mathcal{N}}$. This prove the desire result. ∎
For the dual space of compact operator and nuclear operator, we have the following characterization.
@@
([5, Proposition 16.7]) Let $E$ and $F$ be Banach spaces such that either $E^{\prime\prime}$ or $F^{\prime}$ has the approximation property, and either $E^{\prime\prime}$ or $F^{\prime}$ has the Radon–Nikodým property. Then the following isometric identities hold:
$\displaystyle(\mathcal{K}(E,F))^{\prime}$ $\displaystyle=\mathcal{N}\left(E^{\prime},F^{\prime}\right),$
$\displaystyle(\mathcal{K}(E,F))^{\prime\prime}$ $\displaystyle=\left(\mathcal{N}\left(E^{\prime},F^{\prime}\right)\right)^{\prime}=\mathcal{L}\left(E^{\prime\prime},F^{\prime\prime}\right),$
where the duality pairings are given by
$\displaystyle\langle S,T\rangle_{\mathcal{K},\mathcal{N}}$ $\displaystyle=\sum_{n=1}^{\infty}\left\langle S^{*}y_{n}^{\prime},x_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}},$
$\displaystyle\langle T,S\rangle_{\mathcal{N},\mathcal{L}}$ $\displaystyle=\sum_{n=1}^{\infty}\left\langle y_{n}^{\prime},Sx_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}},$
for $T=\sum_{n=1}^{\infty}x_{n}^{\prime\prime}\otimes y_{n}^{\prime}\in\mathcal{N}\left(E^{\prime},F^{\prime}\right)$.
@@
$\displaystyle\langle S,T\rangle$ $\displaystyle=Tr(ST)=Tr(TS)$
$\displaystyle\langle A,B\rangle$ $\displaystyle=Tr(AB)=Tr(BA)$

### v2

@@
Abstract
We characterize the nuclearity of Toeplitz operators $T_{\mu}:F_{\alpha}^{p}\to F_{\alpha}^{q}$ with Borel measure symbols for $1\leq p,q\leq\infty$. For positive measures $\mu$ and $q\leq p$, we provide necessary and sufficient conditions in terms of the Berezin transform and establish a rigidity property for nuclearity across this range. In the case $p<q$, we obtain separate necessary and sufficient conditions, indicating that the Berezin transform alone is insufficient for a complete characterization. Our results extend to Fock spaces on $\mathbb{C}^{n}$.
keywords:
@@ Abstract
articletype: RESEARCH ARTICLE
{Frontmatter}
1 Introduction
Toeplitz operators on Fock spaces constitute a fundamental class of non-selfadjoint operators in complex analysis and operator theory and have attracted a great deal of interest. The theory of Toeplitz operators on Fock spaces has been extensively developed in the Hilbert space setting. For the classical Fock space $F_{\alpha}^{2}$, Toeplitz operators with bounded or measure-valued symbols have been studied from various perspectives, including boundedness, compactness, and Schatten class membership, see references [MR2610379, MR882716, MR1277446, MR4107813]. In particular, Isralowitz and Zhu [MR2609242] obtained complete characterizations of boundedness, compactness, and Schatten class properties of Toeplitz operators with positive Borel measure symbols.
Beyond the Hilbert space situation, Hu and Lv [MR2819157, MR3248473] systematically investigated Toeplitz operators acting between different Fock spaces $F_{\alpha}^{p}$ and $F_{\alpha}^{q}$ for $0<p,q<\infty$. They characterized bounded and compact Toeplitz operators with positive Borel measure symbols in terms of averaging functions and Berezin-type transforms. A notable feature of their results is a rigidity phenomenon specific to Fock spaces: boundedness (or compactness) from $F_{\alpha}^{p}$ to $F_{\alpha}^{q}$ for some $p\leq q$ automatically implies boundedness (or compactness) for all such $q$. The groundbreaking work of Hu and Lv has stimulated extensive scholarly investigations into Toeplitz and Hankel operators across diverse weighted Fock spaces [MR3392258, MR3157978, MR2934601].
In contrast to the well-developed theory of bounded and compact Toeplitz operators, considerably less is known about Toeplitz operators belonging to finer operator ideals on Fock spaces, especially in the non-Hilbertian case. Among these ideals, the class of nuclear operators plays a distinguished role in Banach space operator theory. The aim of this paper is to study nuclear Toeplitz operators acting between Fock spaces. Our first main result is presented as follows:
Theorem 1.1.
Let $1\leq q\leq p<\infty,1\leq s\leq\infty$, and let $\mu\geq 0$ satisfy (2.1); then the following statements are equivalent
  1. (1)
    $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$;
  2. (2)
    $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$;
  3. (3)
    $\mu(\mathbb{C})<\infty$.
Moreover,
$\|T_{\mu}\|_{\mathcal{N}}\simeq\mu(\mathbb{C}).$ (1.1)
Since nuclear and trace-class operators coincide on Hilbert spaces, the case $p=q=2$ in our results recovers Zhu’s characterization of trace‑class Toeplitz operators in [MR2609242]. Notice that Zhu’s proof relies heavily on Hilbert space structure, including orthogonal decompositions, positivity, and the classical Schatten class theory. In contrast, when $p\neq 2$, the Fock space $F^{p}_{\alpha}$ is a Banach space lacking Hilbert geometry. The standard techniques used in the Hilbert case do not extend directly. Thus, the characterization of nuclear Toeplitz operators between Banach Fock spaces requires a different approach based on operator ideals. We emphasize that our method provides a more constructive approach to approximating $T_{\mu}$ with nuclear operators compared to Zhu’s proof.
Our main results provide a detailed description of nuclear Toeplitz operators in different parameter regimes. When $1\leq q\leq p<\infty$ and $\mu\geq 0$, we obtain necessary and sufficient conditions for $T_{\mu}$ to be nuclear in terms of the integrability of the Berezin transform of $\mu$. As a consequence, we show that nuclearity in this range exhibits a rigidity property analogous to that observed for boundedness: if $T_{\mu}$ is nuclear from $F_{\alpha}^{p}$ to $F_{\alpha}^{q}$ for some $q\leq p$, then it is nuclear for all such $q$.
When $p<q$, the situation becomes more subtle. In this case, we establish separate necessary and sufficient conditions for nuclearity, showing that the Berezin transform alone does not yield a complete characterization. This reflects an inherent asymmetry between the domain and target spaces in the nuclear setting and distinguishes the nuclear theory from previously studied mapping properties of Toeplitz operators on Fock spaces. We will discuss this problem in Chapter 3.
Berger and Coburn [MR1277446] showed that the set of all Toeplitz operators $T_{\varphi}$ with compact support continuous symbol is trace-norm dense in the trace ideal of $F_{\alpha}^{2}$, and is operator-norm dense in $\mathcal{K}(F_{\alpha}^{2})$. Our second main result generalizes the aforementioned finding :
Theorem 1.2.
Let $C$ denote the set of all Toeplitz operators $T_{\varphi}$ with continuous, compactly supported symbols $\varphi$ on $\mathbb{C}$, and let $1<p,q<\infty$. Then:
  1. (1)
    $C$ is dense in $\mathcal{N}(F^{p}_{\alpha},F^{q}_{\alpha})$ with respect to the nuclear norm.
  2. (2)
    $C$ is norm-dense in $\mathcal{K}(F^{p}_{\alpha},F^{q}_{\alpha})$.
The proofs rely on tools from Banach space operator theory, including the duality properties of nuclear operators, approximation and Radon-Nikodym properties of Fock spaces, which we will present in Section 2.
The paper is organized as follows. In Section 2, we recall basic definitions and preliminary results concerning nuclear operators and Fock spaces. Section 3 is devoted to the characterization of nuclear Toeplitz operators between Fock spaces and contains our main results. In Section 4, we present several trace formulas and establish the density theorem.
2 Preliminary
2.1 Fock Space
@@
Let $\alpha>0$ and $p>0$, the notation $L^{p}_{\alpha}$ refers to the space of all Lebesgue measurable functions $f$ on $\mathbb{C}$ for which $f(z)e^{-\alpha|z|^{2}/2}\in L^{p}(\mathbb{C},\mathrm{d}A)$, where $\mathrm{d}A$ is the Lebesgue area measure on $\mathbb{C}$. For $f\in L^{p}_{\alpha}$, its norm is given by
@@
∥f∥_α,p^p=pα2π∫_C|f(z)e^-α2|z|^2|^pdA(z). Correspondingly, for $\alpha>0$ and $p=\infty$, we denote by $L_{\alpha}^{\infty}$ the space of measurable functions $f$ such that
@@
∥f∥_∞,α=ess  sup{|f(z)|e^-α|z|^2/2:z∈C}<∞. The Fock space $F_{\alpha}^{p}$ is defined as the subspace of $L_{\alpha}^{p}$ consisting of entire functions. It is straightforward to verify that $F_{\alpha}^{p}$ becomes a Banach space when $1\leq p\leq\infty$. Let $f_{\alpha}^{\infty}$ be the closed subspace of $F_{\alpha}^{\infty}$ consisting of $f$ satisfying
@@
lim_|z|→∞f(z)e^-α2|z|^2=0. When $p=2$ , the space $F_{\alpha}^{2}$ is a Hilbert space equipped with the inner product
@@
⟨f,g⟩_F_α^2=απ∫_Cf(z)g(z)e^-α|z|^2 dA(z). As a reproducing kernel Hilbert space, its reproducing kernel is $K_{\alpha}(z,w)=e^{\alpha z\overline{w}}$. The normalized reproducing kernel at z is denoted by $k_{z}(w)=e^{\alpha\overline{z}w-\frac{\alpha}{2}|z|^{2}}$ , which is a unit vector in $F_{\alpha}^{p}$ for every $0<p\leq\infty$ . For $1\leq p<\infty$ , let $p^{\prime}$ be the conjugate exponent satisfying $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. We can identify the dual of $F_{\alpha}^{p}$ with $F_{\alpha}^{p^{\prime}}$ , and the dual of $f_{\alpha}^{\infty}$ with $F_{\alpha}^{1}$ , via the integral pairing ⟨f,g⟩_F_α^2=απ∫_Cf(z)g(z)e^-α|z|^2 dA(z), and under this identification, the norms on the dual spaces and the Fock spaces are equivalent (see [MR2934601, Corollary 2.25, Theorem 2.26]).
@@
dλ_α(z)=απe^-α|z|^2 dA(z). Define the integral operator $P_{\alpha}$ on $L_{\alpha}^{p}$ by
P_αf(z)=∫_CK_α(z,w)f(w) dλ_α(w), f∈L_α^p. It is known that $P_{\alpha}:L_{\alpha}^{p}\to F_{\alpha}^{p}$ is bounded for all $p$ , and it acts as the identity on $F_{\alpha}^{p}$ , that is, $P_{\alpha}f=f$ for $f\in F_{\alpha}^{p}$ (see [MR2934601, Corollary 2.22]).
@@
T_φ(f)=P_α(φf), f∈F_α^2. More generally, if $\mu$ is a complex Borel measure on $\mathbb{C}$ , the Toeplitz operator $T_{\mu}$ is given by
T_μ(f)(z)=απ∫_CK_α(z,w)f(w)e^-α|w|^2 dμ(w), z∈C, provided that the integrability condition
$\int_{\mathbb{C}}|K_{\alpha}(z,w)|e^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty$ (2.1)
holds for every $z\in\mathbb{C}$. Let
@@
The set $K$ is dense in $F_{\alpha}^{p}$ for every $1\leq p<\infty$ and also in $f_{\alpha}^{\infty}$. Under assumption (2.1), the operator $T_{\mu}$ is well-defined on $K$. For a measure $\mu$ satisfying (2.1), its Berezin transform is defined as
@@
2.2 Nuclear Operator
@@
for sequences $(x_{j}^{\prime})\subset X^{\prime}$ and $(y_{j})\subset Y$ satisfying $\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y}<\infty.$ For $x^{\prime}\in X^{\prime}$ and $y\in Y$, we denote by $x^{\prime}\otimes y$ the rank-one operator from $X$ to $Y$ defined by x’⊗y(x)=⟨x,x’⟩y, x∈X. The nuclear norm of $T$ is given by
∥T∥_N:=inf∑_i=1^∞∥x_i^′∥_X’∥y_i∥_Y, where the infimum in this definition is taken over all representations of the type (2.4). We write $\mathcal{N}(X,Y)$ for the class of all nuclear operators mapping $X$ into $Y$. Equipped with the norm $\|\cdot\|_{N}$, it becomes a Banach space. In the case where $E$ and $F$ coincide, we write $\mathcal{N}(E)$ for $\mathcal{N}(E,E)$.
A Banach space $E$ is said to have the approximation property if for every compact set $K\subset E$ and every $\varepsilon>0$, there exists a finite-rank operator $L:E\to E$ such that ∥x-Lx∥<ϵfor all x∈K. For a Banach space $E$ with the approximation property, the nuclear trace of an operator $T\in\mathcal{N}(E)$ can be defined by
Tr_E(T)=∑_j=1^∞⟨x_j,x_j^′⟩, where $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes x_{j}$ is any representation as in (2.4). It follows from [MR1744872] Theorem V.1.2, that the value of the trace does not depend on the particular representation chosen. Standard references for the general theory of operator ideals include [MR1209438, MR1342297, MR1744872, MR2300779].
@@
$G(E)=\int_{\Omega}g\,d\mu\quad\text{holds for all }E\in\Sigma,$
we say that $X$ has the Radon-Nikodym property.
@@
For $1<p<\infty$, the Fock space $F_{\alpha}^{p}$ is a Banach space that enjoys both the approximation property and the Radon-Nikodym property.
Proof 1.
The approximation property follows from [MR4107813, Proposition 2.1]. The Radon-Nikodym property holds because $F_{\alpha}^{p}$ is reflexive(see [MR453964, Corollary 3.13]).
@@
$T^{*}(y^{\prime})(x)$ $=y^{\prime}(Tx)$
$=y^{\prime}(\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y_{j})$
$=\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y^{\prime}(y_{j})$
$=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}(y^{\prime})x_{j}^{\prime}(x).$
Then $T^{*}=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}\otimes x_{j}^{\prime}$, and $\|T^{*}\|_{\mathcal{N}}\leq\sum_{j=1}^{\infty}\|y_{j}^{\prime\prime}\|_{Y^{\prime\prime}}\|x_{j}^{\prime}\|_{X^{\prime}}=\sum_{j=1}^{\infty}\|y_{j}\|_{Y}\|x_{j}^{\prime}\|_{X^{\prime}}$. Taking the infimum over all possible representations (2.4), we obtain $\|T^{*}\|_{\mathcal{N}}\leq\|T\|_{\mathcal{N}}$. On the other hand, let $K\in\mathcal{N}(Y^{\prime},X^{\prime})$. Since $Y$ is reflexive, there exist sequences $(a_{j}^{\prime})\in X^{\prime}$ and $(b_{j})\in Y$ such that $K=\sum_{j=1}^{\infty}b_{j}^{\prime\prime}\otimes a_{j}^{\prime}$. Let $L=\sum_{j=1}^{\infty}a_{j}^{\prime}\otimes b_{j}\in\mathcal{N}(X,Y)$. It is easy to check that $L^{*}=K$ and $\|L\|_{\mathcal{N}}\leq\|K\|_{\mathcal{N}}$. This proves the desired result.
For the dual space of compact operators and nuclear operators, we have the following characterization.
@@
([MR1209438, Proposition 16.7]) Let $E$ and $F$ be Banach spaces such that $E^{\prime\prime}$ or $F^{\prime}$ has the approximation property and: $E^{\prime\prime}$ or $F^{\prime}$ has the Radon-Nikodým property. Then
$(\mathcal{K}(E,F))^{\prime}$ $=\mathcal{N}\left(E^{\prime},F^{\prime}\right)$ $\text{isometrically},$
$(\mathcal{K}(E,F))^{\prime\prime}$ $=\left(\mathcal{N}\left(E^{\prime},F^{\prime}\right)\right)^{\prime}=\mathcal{L}\left(E^{\prime\prime},F^{\prime\prime}\right)$ $\text{isometrically}.$
The duality brackets are given by
$\langle S,T\rangle_{\mathcal{K},\mathcal{N}}$ $=\sum_{n=1}^{\infty}\left\langle S^{*}y_{n}^{\prime},x_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}}$
$\langle T,S\rangle_{\mathcal{N},\mathcal{L}}$ $=\sum_{n=1}^{\infty}\left\langle y_{n}^{\prime},Sx_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}}$
if $T=\sum_{n=1}^{\infty}x_{n}^{\prime\prime}\otimes y_{n}^{\prime}\in\mathcal{N}\left(E^{\prime},F^{\prime}\right)$.
@@
$\langle S,T\rangle$ $=\mathrm{Tr}(ST)=\mathrm{Tr}(TS)$
$\langle A,B\rangle$ $=\mathrm{Tr}(AB)=\mathrm{Tr}(BA)$